OpenStudy (anonymous):

Part 1: Find the area of a rectangular object which has a length of 2√3 inches and a width of 3√6 inches. (3 points) Part 2: Explain, in complete sentences, how you arrived at the simplified answer and give the final solution in simplified radical form. (2 points) Part 3: What type of object in your home or school might this be? (1 point)

OpenStudy (anonymous):

\[\left(2 \sqrt{3}\right)*\left( 3 \sqrt{6}\right)=6\sqrt{18}=6*\left(3\sqrt{2}\right)=18 \sqrt{2} \]A piece of paper.

OpenStudy (anonymous):

thank u